# Symmetry classification correspondence between quadratic Lindbladians and their steady states

### v1

@@ Abstract
Classification is a powerful method for understanding and organizing universal properties and is broadly used in physics and other natural sciences. In closed quantum systems, a many-body Hamiltonian is classified by the behavior under time-reversal, particle-hole, and chiral symmetries, leading to the celebrated Altland-Zirnbauer (AZ) tenfold way classification[1, 2, 3]. While conceptually simple, the tenfold way classification determines universal physical properties of non-interacting fermionic systems. For a gapped fermionic system, the AZ symmetry classes and spatial dimension fully determine if non-trivial topological phases could appear[4, 5, 6]. For a disordered system, the symmetry classes determine the level statistics[7, 8], and effective theory on the boundary[9, 10].
Accounting for the great success of the tenfold way classification theory, it is appealing to develop a symmetry classification theory for open quantum systems[11, 12]. A broad class of open quantum systems can be described by the Lindblad master equation on density matrix $\hat{\rho}$:
@@
The right-hand-side of the master equation is regarded as a linear map $\mathscr{L}[\hat{\rho}]$ on the space of density operators, and $\mathscr{L}$ is named a superoperator. $\mathscr{L}[\hat{\rho}]$ plays a role similar to that of the many-body Hamiltonian for closed systems. But it possesses complex eigenvalues in general. The non-Hermiticity of $\mathscr{L}$ forbids a direct application of AZ tenfold way theory. Fortunately, there exists a theory that generalizes the AZ classification theory to the non-Hermitian domain, usually called the Bernard-LeClair (BL) symmetry classification theory[13]. The BL classification theory was recently demonstrated to be a successful framework for classifying non-Hermitian matrices[14, 15, 16, 17, 18]. Several recent works have applied the BL symmetry classification theory to classify Lindbladian superoperators $\mathscr{L}$[19, 20, 21], and unveiled some universal topological[22, 23, 24, 25, 26, 27, 28] and quantum chaos[29, 30, 31, 32, 33, 34, 35, 36, 37, 38] properties of open quantum systems.
@@
Figure 1: Symmetry classification correspondence of $\mathscr{L}$ and $\hat{\rho}_{\text{ss}}$. There are 27 possible BL symmetry classes of $\mathscr{L}$ with particle conservation. Among them 22 classes lead to infinite temperature steady state. The remaining 5 classes are in one-to-one correspondence with 5 AZ classes of steady states. For the definition of each class, see Tab.3.
However, an outstanding question still remains largely unexplored: How the symmetry classification of the Lindbladian affects the steady state, especially the symmetry classification of its steady states? Intuitively, the structure of Lindbladian $\mathscr{L}$ fully determines the steady state $\hat{\rho}_{\text{ss}}$. So it is expected that symmetry classification of $\mathscr{L}$ also determines the physical properties of $\hat{\rho}_{\text{ss}}$. But when it comes to symmetry classification, a division immediately arises: the steady state $\hat{\rho}_{\text{ss}}$ is a Hermitian operator and is therefore subject to AZ instead of the BL classification theory. The classification of $\mathscr{L}$ and $\hat{\rho}_{\text{ss}}$ utilize different frameworks. Their connections remain to be clarified.
@@
In this paper we bridge this gap for quadratic fermionic Lindbladians with U(1) particle conservation. We formulate a concrete correspondence between BL symmetry classes of $\mathscr{L}$ and AZ symmetry classes of $\hat{\rho}_{\text{ss}}$, assuming $\hat{\rho}_{\text{ss}}$ is unique. We extract the matrix representation $L_{\text{eff}}$ of the Lindbladian by defining proper fermionic superoperators, a method dubbed third quantization[39, 40]. Since $\mathscr{L}$ must preserve trace and Hermiticity, $L_{\text{eff}}$ must be a highly structured matrix. We find $L_{\text{eff}}$ has 27 possible non-Hermitian symmetry classes. The structure of $L_{\text{eff}}$ also limits the form of symmetry transformations, and allows us to build connections to the steady state symmetries. Based on this, we further deduce that $\hat{\rho}_{\text{ss}}$ of 22 classes is the maximally mixed state. Each of the remaining 5 classes has a one-to-one correspondence with 5 of the 10 steady state AZ classes. The results are illustrated in Fig.1
@@
Symmetry classification overview. Let us first summarize the notations and set the problem. The focus of this paper is Lindbladians satisfying the following conditions. $\hat{H}=H_{ij}\hat{c}_{i}^{\dagger}\hat{c}_{j}$ is a quadratic operator of fermion operators $\hat{c}^{\dagger}$ and $\hat{c}$. Dissipation operators $\hat{L}_{\mu}$ are linear superpositions of either creation operators $\hat{L}_{\mu}=D^{g}_{\mu i}\hat{c}_{i}^{\dagger}$ or annihilation operators $\hat{L}_{\mu}=D^{l}_{\mu i}\hat{c}_{i}$. The total number of fermion modes is $N$. Thus $H$ is a $N\times N$ matrix. We also introduce $(M_{l})_{ij}=D^{l*}_{\mu i}D^{l}_{\mu j}$ and $(M_{g})_{ij}=D^{g*}_{\mu i}D^{g}_{\mu j}$ to characterize the structure of loss and gain. We note both $M_{l}$ and $M_{g}$ are positive semi-definite.
@@
Recall that, in the classification of topological insulators, the AZ classification theory is applied to the matrix representation $H$ of a quadratic Hamiltonian $\hat{H}$. For quadratic Lindbladians $\mathscr{L}$, as pointed out in, we can extract a similar matrix representation. A quadratic Lindbladian can be written as a bilinear form using a set of fermionic superoperators, which we denote as $\mathscr{a}$ and $\mathscr{b}$, or more succinctly, super-fermions (SFs),
@@
Physically, $\mathscr{a}$ and $\mathscr{b}$ represent the fermions in the left and right subspace after mapping the density matrix to a quantum state via Choi-Jamiolkowski isomorphism[41, 42].
With SFs, we are able to write $\mathscr{L}$ as a quadratic $\mathscr{L}$ has two subspaces, $\mathscr{L}_{\pm}$, defined by $\mathcal{P}^{F}=\pm 1$. Here $L_{\text{eff}}$ is only the representation of $\mathscr{L}_{+}$. On one hand, only the $\mathscr{L}_{+}$ subspace has steady states. On the other hand, the matrix representation of $\mathscr{L}_{-}$ is $L_{\text{eff}}^{\prime}=(\sigma_{z}\otimes\mathbb{I}_{N})L_{\text{eff}}(\sigma_{z}\otimes\mathbb{I}_{N})^{\dagger}$. $L_{\text{eff}}^{\prime}$ has the same symmetry class as $L_{\text{eff}}$. Therefore, we simply call $\mathscr{L}_{+}$ the full Lindbladian $\mathscr{L}$. It does not influence the subsequent discussion of symmetry classification.,
@@
Here $\mathscr{a}$ is the shorthand for the column vector $(\mathscr{a}_{1},\cdots,\mathscr{a}_{N})^{T}$, and the same notation is used for $\mathscr{b}$. The constant term $\text{Tr}(\cdots)$ guarantees that the real part of all the eigenvalues of $\mathscr{L}$ is non-positive. We denote the large $2N\times 2N$ matrix representation as $L_{\text{eff}}$. In this way, we translate the classification of superoperators $\mathscr{L}$ to non-Hermitian matrices $L_{\text{eff}}$ [^2] In previous studies, e.g.[19, 25], the damping matrix $X=iH^{T}-M_{l}^{T}-M_{g}$ was subject to classification. While $X$ contains all the spectrum information, it does not contain the eigenstate information of $\mathscr{L}$. To address the relation between $\mathscr{L}$ and $\hat{\rho}_{\text{ss}}$, here we classify $L_{\text{eff}}$, which contains both spectrum and eigenstate information..
@@
$U_{K,C,Q,P}$ are unitary matrices. $\epsilon_{K,C,Q}=\pm 1$, $\eta_{K,C}=\pm 1$ denote the sign of the symmetries. If more than one symmetry is present, symbols such as $\epsilon_{PQ}=\pm 1$ are introduced to represent the commutation or anti-commutation relations between the symmetries. We denote the Q symmetry with $\epsilon_{Q}=\pm 1$ as the Q± symmetry, and use similar notations for other symmetries. Counting the presence and sign of symmetries, there are 54 possible symmetry classes [^3] The number of possible symmetry classes of a non-Hermitian matrix depends on the type of complex-energy gap. For Lindbladians, the dissipation gap is defined as the separation of eigenvalue’s real part from the steady state. So the line gap scenario is proper, where 54 symmetry classes exist in total[17]..
@@
The goal of this work is to find the relation between BL symmetry classification of $L_{\text{eff}}$ and AZ symmetry classification of $G$.
@@
Summary of main results. Before going to the detailed analysis, we summarize the main results of this work.
(1) From 54 to 27: A Lindbladian operator must retain the Hermiticity-preserving property. That is, when $\mathscr{L}$ operates on any $\hat{\rho}$, $\mathscr{L}[\hat{\rho}]$ should always be a hermitian operator. This constraint is equivalent to a Q- symmetry $(\sigma_{y}\otimes\mathbb{I}_{N})L_{\text{eff}}(\sigma_{y}\otimes\mathbb{I}_{N})^{\dagger}=-L_{\text{eff}}^{\dagger}$. It eliminates 27 symmetry classes.
(2) From 27 to 5: Among the remaining 27 classes, there are 22 symmetry classes in which loss and gain terms share the same structure, leading to an infinite-temperature structureless steady state. Only Lindbladians in five symmetry classes can have a nontrivial steady state.
@@
A full list can be found in Tab.3 within the End Matter (EM).
Derivation sketch. Here we sketch the main ideas to derive the symmetry classification correspondence. The main difficulties in building connections between $L_{\text{eff}}$ and $G$ are: (1) Although $L_{\text{eff}}$ determines $G$, the concrete functional relation is complicated; (2) Symmetry transformations of $L_{\text{eff}}$ and $G$ have different dimensions in their matrix representation. To overcome these difficulties, a crucial observation is that only $H$, $M_{l}$, and $M_{g}$ are free parameters. They determine both $L_{\text{eff}}$ and $G$. As a result, we seek to reduce both the symmetry classification of $L_{\text{eff}}$ and $G$ to constraints on $H$, $M_{l}$ and $M_{g}$.
For $G$, we introduce the correlation matrix $(C_{\text{ss}})_{ij}=\text{Tr}(\hat{\rho}_{\text{ss}}\hat{c}^{\dagger}_{i}\hat{c}_{j})$ to express $\hat{\rho}_{\text{ss}}$, since it is guaranteed to be Gaussian. $C_{\text{ss}}$ is related to $G$ by $C_{\text{ss}}=1/(e^{G^{T}}+1)$. So symmetries of $G$ can be equivalently expressed as constraints of $C_{\text{ss}}$,
@@
where $X=iH^{T}-M_{l}^{T}-M_{g}$. When the equation has a unique solution (unique steady state), constraints of $C_{\text{ss}}$ can be equivalently expressed as constraints on $H$, $M_{l}$ and $M_{g}$.
For $L_{\text{eff}}$, it is crucial to observe that $L_{\text{eff}}$ is a highly structured matrix, which puts constraints on the structure of its symmetry transformation $U$. For generic parameters, we find $U$ takes the form $U=\sigma\otimes V$, where $V$ is a $N\times N$ unitary and $\sigma$ is a Pauli or identity matrix. The concrete form of $\sigma$ varies among symmetries. The form of $U$ has a clear physical meaning. $V$ accounts for transformations inside each of the spaces of $\mathscr{a}$ and $\mathscr{b}$. $\sigma$ represents possible swaps between them. Since $\mathscr{a}_{i}$ and $\mathscr{b}_{i}$ are actually the same fermion mode $\hat{c}_{i}$ acting on different sides of density matrix, they should be transformed consistently by $V$. As a result, the symmetry transformation is reduced from $U$ to $V$, which puts constraints on $H$, $M_{l}$ and $M_{g}$.
Below we first state and illustrate the reduction of symmetry transformation. Then, we will perform a BL classification of $L_{\text{eff}}$ and, with the help of these results, discuss the effects on steady state.
@@
Symmetry transformation reduction. The results are summarized in Tab.1. Here we consider only P and C± symmetries, for a reason that will be clear later. The second column of the table shows the form of symmetry transformation $U$ of each symmetry. The third column shows the constraints on matrices $H$, $M_{l}$, and $M_{g}$. A comprehensive derivation can be found in the EM. To illustrate the validity, we show two examples below
$=-L_{\text{eff}}.$ (9)
This gives us $VM_{l}V^{\dagger}=-M_{g}^{T}$ by considering the $(1,2)$ off-diagonal block, contradicting the positive semi-definiteness of $M_{l}$ and $M_{g}$. Here we see the form of $L_{\text{eff}}$ and positive semi-definiteness of $M_{l,g}$ result in the constraint on symmetry transformation.
@@
Classification of $L_{\text{eff}}$ and steady state properties Now we analyze the BL symmetry classification of $L_{\text{eff}}$, and discuss the steady state properties. We assume that $\mathscr{L}$ has a unique steady state. The full results are summarized comprehensively in Tab.3.
First of all, as previously mentioned, $L_{\text{eff}}$ has an inherent Q- symmetry expressed as $(\sigma_{y}\otimes\mathbb{I}_{N})L_{\text{eff}}(\sigma_{y}\otimes\mathbb{I}_{N})^{\dagger}=-L_{\text{eff}}^{\dagger}$. There are only 27 possible symmetry classes compatible with this condition, which are therefore suitable for $L_{\text{eff}}$. With the Q- symmetry present, K and C symmetries become indistinguishable. So only P and C± symmetries need to be considered for subsequent classification. That is the reason why we only consider these symmetries before.

### v2

@@ Abstract
Symmetry plays a significant role in understanding universal physical properties of quantum many-body systems. Many-body Hamiltonians are classified by the behavior under time-reversal, particle-hole, and chiral symmetries, leading to the celebrated Altland-Zirnbauer (AZ) tenfold way classification[1, 2, 3]. The AZ classification provides a powerful framework to understanding fermionic systems. This classification dictates the appearance of topological phases in translational-invariant gapped fermions based on symmetry class and dimensionality[4, 5, 6]. Furthermore, in chaotic systems, the symmetry class determines universal random matrix correlations, defining level statistics[7, 8] and further the effective theory[9, 10].
Recent studies of symmetry classification extend to open quantum systems, typically described by the Lindblad master equation[11, 12]
@@
For such systems, both the long-time stationary state and the transient dynamical phenomenon are of fundamental importance. The former is described by the steady state density matrix $\hat{\rho}_{\text{ss}}$. The symmetry and topology of $\hat{\rho}_{\text{ss}}$, particularly for fermionic Gaussian states amenable to AZ tenfold way classification, have been extensively investigated[13, 14, 15, 16, 17, 18, 19, 20]. Transient phenomena, on the other hand, is described by the dynamical generator $\mathscr{L}$, dubbed the Lindbladian superoperator. When regarded as a linear map in the operator space, $\mathscr{L}$ is in general non-Hermitian, preventing a direct AZ classification. Fortunately, the Bernard-LeClair (BL) classification for non-Hermitian matrices[21], recently well-developed and validated [22, 23, 24, 25, 26], offers a suitable framework. Applying the BL classification theory to $\mathscr{L}$, several recent work unveiled universal topological[27, 28, 29, 30, 31, 32, 33] and quantum chaos[34, 35, 36, 37, 38, 39, 40, 41, 42, 43] properties of open quantum systems.
@@
Figure 1: Symmetry classification correspondence between $\mathscr{L}$ and $\hat{\rho}_{\text{ss}}$. Of 27 BL symmetry classes for $\mathscr{L}$, 22 yield an infinite temperature steady state. The remaining 5 classes map one-to-one to 5 steady state AZ classes. For definition of each class, see Tab.3.
However, a fundamental question still remains largely unexplored: How the symmetry classification of the Lindbladian affects the steady state, especially the symmetry classification of its steady state? Given the physical intuition that transient dynamics gradually shape the steady state, we expect the symmetry classification of $\mathscr{L}$ to dictate the physical properties of $\hat{\rho}_{\text{ss}}$. However, a gap arises when considering their symmetry classifications: $\mathscr{L}$ and $\hat{\rho}_{\text{ss}}$ are described within different theoretical frameworks. Furthermore, the BL classification theory for $\mathscr{L}$ comprises 54 possible symmetry classes[22, 23, 24, 25, 26], while the AZ classification for $\hat{\rho}_{\text{ss}}$ encompasses only ten. Due to this divergence, the connections between symmetry classifications of $\mathscr{L}$ and $\hat{\rho}_{\text{ss}}$ are yet to be clear.
@@
This paper bridge the gap for quadratic fermionic Lindbladians with U(1) particle conservation. We formulate a precise correspondence between the BL symmetry classes of $\mathscr{L}$ and the AZ symmetry classes of its unique steady state $\hat{\rho}_{\text{ss}}$. We extract the matrix representation $L_{\text{eff}}$ of the Lindbladian by defining proper fermionic superoperators, a method dubbed third quantization[44, 45]. Since $\mathscr{L}$ must preserve trace and Hermiticity, $L_{\text{eff}}$ is a highly structured matrix, allowing for 27 distinct BL symmetry classes. The structure of $L_{\text{eff}}$ also constraints the form of symmetry transformations, and allows us to build connections to the steady state symmetries. Consequently, we further deduce that $\hat{\rho}_{\text{ss}}$ of 22 classes is the maximally mixed state. Each of the remaining 5 classes has a one-to-one correspondence with 5 of the 10 steady state AZ classes. The results are illustrated in Fig.1
I Symmetry classification overview.
@@
The focus of this paper is Lindbladians satisfying the following conditions. $\hat{H}=H_{ij}\hat{c}_{i}^{\dagger}\hat{c}_{j}$ is a quadratic operator of fermion operators $\hat{c}^{\dagger}$ and $\hat{c}$. Dissipation operators $\hat{L}_{\mu}$ are linear superpositions of either creation operators $\hat{L}_{\mu}=D^{g}_{\mu i}\hat{c}_{i}^{\dagger}$ or annihilation operators $\hat{L}_{\mu}=D^{l}_{\mu i}\hat{c}_{i}$. The total number of fermion modes is $N$. Thus $H$ is a $N\times N$ matrix. We also introduce $(M_{l})_{ij}=D^{l*}_{\mu i}D^{l}_{\mu j}$ and $(M_{g})_{ij}=D^{g*}_{\mu i}D^{g}_{\mu j}$ to characterize the structure of loss and gain. We note both $M_{l}$ and $M_{g}$ are positive semi-definite.
@@
In the classification of topological insulators, the AZ classification theory is applied to the matrix representation $H$ of a quadratic Hamiltonian $\hat{H}$. For quadratic Lindbladians $\mathscr{L}$, as pointed out in[44, 45], we can extract a similar matrix representation. A quadratic Lindbladian can be written as a bilinear form using a set of fermionic superoperators, which we denote as $\mathscr{a}$ and $\mathscr{b}$, or more succinctly, super-fermions (SFs),
@@
Physically, $\mathscr{a}$ and $\mathscr{b}$ are the fermions in the left and right subspaces after mapping the density matrix to a quantum state via Choi-Jamiolkowski isomorphism[46, 47].
With SFs, we are able to write $\mathscr{L}$ as a quadratic superoperator
@@
Here $\mathscr{a}$ is the shorthand for the column vector $(\mathscr{a}_{1},\cdots,\mathscr{a}_{N})^{T}$, and the same for $\mathscr{b}$. The constant term $\text{Tr}(\cdots)$ guarantees that the real parts of all the eigenvalues of $\mathscr{L}$ are non-positive. We denote the large $2N\times 2N$ matrix representation as $L_{\text{eff}}$. In this way, we translate the classification of superoperators $\mathscr{L}$ to non-Hermitian matrix $L_{\text{eff}}$.
@@
$U_{K,C,Q,P}$ are unitary matrices. $\epsilon_{K,C,Q}=\pm 1$, $\eta_{K,C}=\pm 1$ denote the sign of the symmetries. If more than one symmetry is present, symbols like $\epsilon_{PQ}=\pm 1$ are introduced to represent the commutation or anti-commutation relations between the symmetries. We denote Q symmetry with $\epsilon_{Q}=\pm 1$ as the Q± symmetry, and other symmetries analogously. Counting the presence and sign of symmetries, there are 54 possible symmetry classes.
@@
The goal of this work is to perform BL classification on $L_{\text{eff}}$ and find its relation to AZ classification of $G$.
@@
II Summary of main results.
We summarize the main results of this work as follows.
(1) From 54 to 27: A Lindbladian operator must retain the Hermiticity-preserving property. When $\mathscr{L}$ operates on any $\hat{\rho}$, $\mathscr{L}[\hat{\rho}$] should always be a hermitian operator. This constraint is equivalent to a Q- symmetry $(\sigma_{y}\otimes\mathbb{I}_{N})L_{\text{eff}}(\sigma_{y}\otimes\mathbb{I}_{N})^{\dagger}=-L_{\text{eff}}^{\dagger}$. It eliminates 27 symmetry classes.
(2) From 27 to 5: Among the remaining 27 classes, there are 22 classes in which loss and gain terms share the same structure, leading to an infinite-temperature structureless steady state. Only Lindbladians in five symmetry classes can have a nontrivial steady state.
@@
A full list can be found in Tab.3 within Appendix. Below we present the derivation.
III Derivation sketch.
Building connections between $L_{\text{eff}}$ and $G$ presents two main difficulties: (1) Although $L_{\text{eff}}$ determines $G$, the concrete relation is complicated; (2) Symmetry transformations of $L_{\text{eff}}$ and $G$ have different dimensions in their matrix representation. To overcome these, a crucial observation is that only $H$, $M_{l}$, and $M_{g}$ are free parameters determining both $L_{\text{eff}}$ and $G$. As a result, our strategy is to reduce both symmetry classifications of $L_{\text{eff}}$ and $G$ to constraints on $H$, $M_{l}$ and $M_{g}$.
For $G$, we introduce the correlation matrix $(C_{\text{ss}})_{ij}=\text{Tr}(\hat{\rho}_{\text{ss}}\hat{c}^{\dagger}_{i}\hat{c}_{j})$ to express $\hat{\rho}_{\text{ss}}$, since it is guaranteed to be Gaussian. $C_{\text{ss}}$ is related to $G$ by $C_{\text{ss}}=1/(e^{G^{T}}+1)$. Symmetries of $G$ translate directly to constraints of $C_{\text{ss}}$,
@@
where $X=iH^{T}-M_{l}^{T}-M_{g}$. When the equation has a unique solution (unique steady state), constraints of $C_{\text{ss}}$ are further translated to constraints on $H$, $M_{l}$ and $M_{g}$.
For $L_{\text{eff}}$, it is crucial to observe that the highly structured form of $L_{\text{eff}}$ constrains its symmetry transformation $U$. For generic parameters, we find $U$ takes the form $U=\sigma\otimes V$, where $V$ is a $N\times N$ unitary and $\sigma$ is a Pauli or identity matrix. The concrete form of $\sigma$ varies among symmetries. The form of $U$ has a clear physical meaning. $V$ accounts for transformations inside each of the spaces of $\mathscr{a}$ and $\mathscr{b}$. $\sigma$ represents possible swaps between them. Since $\mathscr{a}_{i}$ and $\mathscr{b}_{i}$ are actually the same fermion mode $\hat{c}_{i}$ acting on different sides of density matrix, they should be transformed consistently by $V$. As a result, the symmetry transformation is reduced from $U$ to $V$, which can be further translated to constraints on $H$, $M_{l}$ and $M_{g}$.
IV Symmetry transformation reduction.
The constraints on $H$, $M_{l}$, and $M_{g}$ deduced from each symmetry are summarized in Tab.1. Here we consider only P and C± symmetries, for a reason that will be clear later. The second column of the table shows the form of symmetry transformation $U$ of each symmetry. The third column shows the constraints on matrices $H$, $M_{l}$, and $M_{g}$. A comprehensive derivation can be found in the Supplementary Materials. To illustrate the validity, we show two examples below
$=-L_{\text{eff}}.$ (9)
This gives us $VM_{l}V^{\dagger}=-M_{g}^{T}$ by considering the $(1,2)$ off-diagonal block, contradicting the positive semi-definiteness of $M_{l}$ and $M_{g}$. It’s clear that what matter are the form of $L_{\text{eff}}$ and positive semi-definiteness of $M_{l,g}$.
@@
V Classification of $L_{\text{eff}}$ and steady state properties
Now we analyze the BL symmetry classification of $L_{\text{eff}}$, and discuss the steady state properties. We assume that $\mathscr{L}$ has a unique steady state. The full results are summarized comprehensively in Tab.3.
First of all, as previously mentioned, $L_{\text{eff}}$ has an inherent Q- symmetry expressed as $(\sigma_{y}\otimes\mathbb{I}_{N})L_{\text{eff}}(\sigma_{y}\otimes\mathbb{I}_{N})^{\dagger}=-L_{\text{eff}}^{\dagger}$. There are only 27 symmetry classes including this symmetry, which are therefore suitable for $L_{\text{eff}}$. With the Q- symmetry present, K and C symmetries become indistinguishable. So only P and C± symmetries need to be considered for subsequent classification. That is the reason why we only consider them before.